How long is each side of an octagon that is 16 ft. across?
Answer by Alan3354(69443)

How long is each side of an octagon that is 16 ft. across?
---------------
If it's 16 ft across the flats:
The central angles are 360/8 = 45 degs
The apothem = 16/2 = 8 ft
tan(45/2) = (s/2)/8
s = 2*8*tan(22.5)
s =~ 6.627 feet